XluisX:

Translational motion x = x 0 + v 0 t + 1/2at2

justus:

v(t) = v0+a0t (2) a(t) = a0 (3) This is motion with constant acceleration (a(t) = constant).

XluisX:

What formula did you use was it the formula to calculate speed?

justus:

For constant acceleration

XluisX:

For the equation itself

XluisX:

It has to do with some formula but idk which one to use

XluisX:

i hate physics :/

justus:

one moment

justus:

I understand how you feel

XluisX:

Stressed

justus:

what's the question that it's giving you about this formula?

justus:

this formula means distance=initial_distance+velocity*time +1/2acceleration*t^2

XluisX:

Holup

justus:

okay

XluisX:

1 attachment
justus:

If you have all of these variables, such as x0-initial displacement, v0-initial velocity, t-time, and a-acceleration, you may plug them in to determine the displacement calculated in the question. Meters, kilometers, etc.  You will get displacement if you add in the variables.